# Improving Nonpreemptive Multiserver Job Scheduling with Quickswap

### v1

@@ 5.1 Throughput-optimality of MSFQ
We specifically use the continuous-time Foster-Lyapunov theorem [36]. We must demonstrate that this Lyapunov function has three properties, two of which are defined with reference to the drift $E[G\circ V(\cdot)]$, where $G$ is the instantaneous generator operator of the system.
@@
    There exists a finite set of states $B$ and a positive constant $\delta>0$ such that for all states outside of $B$, $E[G\circ V(\cdot)]\leq-\delta$.
@@
    There exists a constant $C$ such that for all states, $E[G\circ V(\cdot)]\leq C$.
@@
Demonstrating Item 1 is the primary challenge. Item 3 holds with $D=0$. Item 2 follows from the fact that $V$ is linear in the two unbounded inputs $n_{1},n_{k}$, which change by at most 1 upon any transition. Furthermore, from any state, the total transition rate is at most $\lambda+\max\{k\mu_{1},\mu_{k}\}$. Hence, there exists an upper bound on drift from any system state.
It thus remains to prove Item 1, which we show holds with $\delta=\epsilon/2$ and the following finite exception set $B$, consisting of all states in phase 2 in which $n_{1}=k$, so phase 2 has the potential to end, and where $n_{k}$ is below a threshold, as well as the empty state:
@@
We specialize our argument based on the current phase of the system $(z=1,2,3,4)$, and whether the system is a state on the border of switching phases.
@@
$\displaystyle E[G\circ V(\cdot)]=\frac{\lambda_{1}}{k\mu_{1}}+\frac{\lambda_{k}}{\mu_{k}}-u_{k}=\frac{\lambda_{1}}{k\mu_{1}}+\frac{\lambda_{k}}{\mu_{k}}-1=-\epsilon\leq-\delta,$
@@
Next, consider states in phase 1 that are on the border of switching to phase 2. In the subsequent phase 2 state, $n_{k}=0$, because the MSFQ policy only switches from serving large jobs to small jobs when there are no large jobs remaining. As a result, the $\mathbbm{1}\{z=2\}$ term in (1) is 0 when beginning phase 2, so Item 1 holds throughout phase 1.
@@
$\displaystyle E[G\circ V(\cdot)]=\frac{\lambda_{1}}{k\mu_{1}}-\frac{u_{1}}{k}+\frac{\lambda_{k}}{\mu_{k}}+\frac{\epsilon\lambda_{k}}{2\lambda_{k}}=-\epsilon+\frac{\epsilon}{2}=-\epsilon/2=-\delta.$
@@
First, in the case where $u_{1}=k$, the servers are fully occupied by small jobs. In this case, if $n_{1}>k$, the $\mathbbm{1}\{z=2\}$ term in (1) does not change on the next arrival or completion, because $u_{1}=k$ will remain true after the next event. Thus,
$\displaystyle E[G\circ V(\cdot)]=\frac{\lambda_{1}}{k\mu_{1}}-\frac{u_{1}}{k}+\frac{\lambda_{k}}{\mu_{k}}=-\epsilon\leq-\delta.$
If $u_{1}=k$ and $n_{1}=k$, the indicator term has a negative instantaneous drift: If a job completes, $n_{1}=n-1$, so the indicator term becomes $c^{k}-1$, while under any other event, the indicator term remains $c^{k}$. Thus,
$\displaystyle E[G\circ V(\cdot)]=\frac{\lambda_{1}}{k\mu_{1}}-\frac{u_{1}}{k}+\frac{\lambda_{k}}{\mu_{k}}-k\mu_{1}=-\epsilon-k\mu_{1}\leq-\delta.$
In the case where $\ell<u_{1}<k$, some servers are idle, and the system is in phase 3, so small jobs continue to enter service. In this case, we upper bound the drift of $V$ as follows:
$\displaystyle E[G\circ V(\cdot)]$ $\displaystyle=\frac{\lambda_{1}}{k\mu_{1}}-\frac{u_{1}}{k}+\frac{\lambda_{k}}{\mu_{k}}+\lambda_{1}(c^{k-u_{1}}-c^{k-(u_{1}+1)})+u_{1}\mu_{1}(c^{k-u_{1}}-c^{k-(u_{1}-1)})$
@@
$\displaystyle E[G\circ V(\cdot)]\leq 1+c^{k-u_{1}}(1-c)(\lambda_{1}c^{-1}-(\ell-1)\mu_{1})\leq 1+2^{1}(-1)(1)=-1\leq-\epsilon\leq-\delta\,.$
In the case where $0<u_{1}\leq\ell$: the system is in phase 4, and small jobs are blocked from entering service. In this case, arriving jobs do not increase $u_{1}$, so the drift is much simpler:
$\displaystyle E[G\circ V(\cdot)]$ $\displaystyle=\frac{\lambda_{1}}{k\mu_{1}}-\frac{u_{1}}{k}+\frac{\lambda_{k}}{\mu_{k}}+u_{1}\mu_{1}(c^{k-u_{1}}-c^{k-(u_{1}-1)})$
@@
Recalling that $c\geq 2$ and that $c\geq\frac{1}{\mu}$, we find that $E[G\circ V]\leq-1$, which completes this step.
@@
The remaining case within phases 3 and 4 is the case where $u_{1}=0$. In this case, the system must be empty, because we would otherwise switch to phase 1. Recall that this case is in the finite set $B$ of exceptional high-drift states, so it does not affect Item 1.
@@
Having handled states in phases 3 and 4 that do not transition directly to phase 1, we now verify the drift condition for states where the system may transition from phase 3 or 4 to phase 1. When entering phase 1, $u_{k}=1$, so $u_{1}=0$: All small jobs in service have been completed. As a result, the $\mathbbm{1}\{z\not\in\{1,2\}\}$ term in (1) is 0, as desired.

### v2

@@ Acknowledgements
We specifically use the continuous-time Foster-Lyapunov theorem [tweedie_sufficient_1975]. We must demonstrate that this Lyapunov function has three properties, two of which are defined with reference to the drift $\mathbb{E}[G\circ V(\cdot)]$, where $G$ is the instantaneous generator operator of the system.
@@
    There exists a finite set of states $B$ and a positive constant $\delta>0$ such that for all states outside of $B$, $\mathbb{E}[G\circ V(\cdot)]\leq-\delta$.
@@
    There exists a constant $C$ such that for all states, $\mathbb{E}[G\circ V(\cdot)]\leq C$.
@@
Demonstrating Property 1 is the primary challenge. Property 3 holds with $D=0$. Property 2 follows from the fact that $V$ is linear in the two unbounded inputs $n_{1},n_{k}$, which change by at most 1 upon any transition. Furthermore, from any state, the total transition rate is at most $\lambda+\max\{k\mu_{1},\mu_{k}\}$. Hence, there exists an upper bound on drift from any system state.
It thus remains to prove Property 1, which we show holds with $\delta=\epsilon/2$ and the following finite exception set $B$, consisting of all states in phase 2 in which $n_{1}=k$, so phase 2 has the potential to end, and where $n_{k}$ is below a threshold, as well as the empty state:
@@
We specialize our argument based on the current phase of the system $(z=1,2,3,4)$, and whether the system is in a state on the border of switching phases.
@@
$\displaystyle\mathbb{E}[G\circ V(\cdot)]=\frac{\lambda_{1}}{k\mu_{1}}+\frac{\lambda_{k}}{\mu_{k}}-u_{k}=\frac{\lambda_{1}}{k\mu_{1}}+\frac{\lambda_{k}}{\mu_{k}}-1=-\epsilon\leq-\delta,$
@@
Next, consider states in phase 1 that are on the border of switching to phase 2. In the subsequent phase 2 state, $n_{k}=0$, because the MSFQ policy only switches from serving heavy jobs to light jobs when there are no heavy jobs remaining. As a result, the $\mathbbm{1}\{z=2\}$ term in (6) is 0 when beginning phase 2, so Item 1 holds throughout phase 1.
@@
$\displaystyle\mathbb{E}[G\circ V(\cdot)]=\frac{\lambda_{1}}{k\mu_{1}}-\frac{u_{1}}{k}+\frac{\lambda_{k}}{\mu_{k}}+\frac{\epsilon\lambda_{k}}{2\lambda_{k}}=-\epsilon+\frac{\epsilon}{2}=-\epsilon/2=-\delta.$
@@
First, in the case where $u_{1}=k$, the servers are fully occupied by light jobs. In this case, if $n_{1}>k$, the $\mathbbm{1}\{z=2\}$ term in (6) does not change on the next arrival or completion, because $u_{1}=k$ will remain true after the next event. Thus,
$\displaystyle\mathbb{E}[G\circ V(\cdot)]=\frac{\lambda_{1}}{k\mu_{1}}-\frac{u_{1}}{k}+\frac{\lambda_{k}}{\mu_{k}}=-\epsilon\leq-\delta.$
If $u_{1}=k$ and $n_{1}=k$, the indicator term has a negative instantaneous drift: If a job completes, $n_{1}=k-1$, so the indicator term becomes $c^{k}-1$, while under any other event, the indicator term remains $c^{k}$. Thus,
$\displaystyle\mathbb{E}[G\circ V(\cdot)]=\frac{\lambda_{1}}{k\mu_{1}}-\frac{u_{1}}{k}+\frac{\lambda_{k}}{\mu_{k}}-k\mu_{1}=-\epsilon-k\mu_{1}\leq-\delta.$
In the case where $\ell<u_{1}<k$, some servers are idle, and the system is in phase 3, so light jobs continue to enter service. In this case, we upper bound the drift of $V$ as follows:
$\displaystyle\mathbb{E}[G\circ V(\cdot)]$ $\displaystyle=\frac{\lambda_{1}}{k\mu_{1}}-\frac{u_{1}}{k}+\frac{\lambda_{k}}{\mu_{k}}+\lambda_{1}(c^{k-u_{1}}-c^{k-(u_{1}+1)})+u_{1}\mu_{1}(c^{k-u_{1}}-c^{k-(u_{1}-1)})$
@@
$\displaystyle\mathbb{E}[G\circ V(\cdot)]\leq 1+c^{k-u_{1}}(1-c)(\lambda_{1}c^{-1}-(\ell-1)\mu_{1})\leq 1+2^{1}(-1)(1)=-1\leq-\epsilon\leq-\delta\,.$
In the case where $0<u_{1}\leq\ell$: the system is in phase 4, and light jobs are blocked from entering service. In this case, arriving jobs do not increase $u_{1}$, so the drift is much simpler:
$\displaystyle\mathbb{E}[G\circ V(\cdot)]$ $\displaystyle=\frac{\lambda_{1}}{k\mu_{1}}-\frac{u_{1}}{k}+\frac{\lambda_{k}}{\mu_{k}}+u_{1}\mu_{1}(c^{k-u_{1}}-c^{k-(u_{1}-1)})$
@@
Recalling that $c\geq 2$ and that $c\geq\frac{1}{\mu}$, we find that $\mathbb{E}[G\circ V]\leq-1$, which completes this step.
@@
The remaining case within phases 3 and 4 is the case where $u_{1}=0$. In this case, the system must be empty, because we would otherwise switch to phase 1. Recall that this case is in the finite set $B$ of exceptional high-drift states, so it does not affect Property 1.
@@
Having handled states in phases 3 and 4 that do not transition directly to phase 1, we now verify the drift condition for states where the system may transition from phase 3 or 4 to phase 1. When entering phase 1, $u_{k}=1$, so $u_{1}=0$: All light jobs in service have been completed. As a result, the $\mathbbm{1}\{z\not\in\{1,2\}\}$ term in (6) is 0, as desired.